# Out-of-Distribution Detection with a Single Unconditional Diffusion Model

### v1

@@ Abstract
Out-of-distribution (OOD) detection is a critical task in machine learning that seeks to identify abnormal samples. Traditionally, unsupervised methods utilize a deep generative model for OOD detection. However, such approaches necessitate a different model when evaluating abnormality against a new distribution. With the emergence of foundational generative models, this paper explores whether a single generalist model can also perform OOD detection across diverse tasks. To that end, we introduce our method, Diffusion Paths, (DiffPath) in this work. DiffPath proposes to utilize a single diffusion model originally trained to perform unconditional generation for OOD detection. Specifically, we introduce a novel technique of measuring the rate-of-change and curvature of the diffusion paths connecting samples to the standard normal. Extensive experiments show that with a single model, DiffPath outperforms prior work on a variety of OOD tasks involving different distributions. Our code is publicly available at https://github.com/clear-nus/diffpath.
@@ 1 Introduction
Out-of-distribution (OOD) detection, or anomaly detection, seeks to detect abnormal samples that are drawn far from a given distribution. This is a vital problem as deep neural networks are known to be overconfident when making incorrect predictions on OOD samples [1, 2], leading to potential issues in safety-critical applications such as healthcare, finance and criminal justice [3]. Traditionally, OOD detection using only unlabeled data relies on training a generative model on in-distribution (ID) data. Thereafter, measures such as model likelihood or its variants are used as an OOD detection score [4, 5, 6]. An alternative approach is to utilize the excellent sampling capabilities of diffusion models (DMs) to reconstruct corrupted samples, and use the reconstruction loss as an OOD measure [7, 8, 9].
However, these conventional methods necessitate separate generative models tailored to specific ID datasets and require retraining if the ID data evolves, such as in continual learning setups. This prompts the question: can OOD detection be performed using a single generative model? We answer in the affirmative and present Diffusion Paths (DiffPath) in this paper. While the use of a single model for OOD detection has been proposed in the discriminative setting [10], to the best of our knowledge we are the first to explore this for generative models. We believe that the generative setting is particularly salient in light of recent trends where large foundational generative models achieve state-of-the-art (SOTA) results across various tasks [11, 12].
Our method utilizes a single DM trained on a diverse dataset (e.g., ImageNet). In a departure from prior works that utilize variants of likelihoods [6, 5, 4] or reconstruction losses [7, 8, 9], we propose to perform OOD detection by measuring characteristics of the forward diffusion trajectory, specifically its rate-of-change and curvature, which can be computed from the score predicted by the diffusion model. We provide theoretical and empirical analyses that motivate these quantities as useful OOD detectors; their magnitudes are similar for samples from the same distribution and different otherwise. We summarize our contributions as follows:
Figure 1: Illustration of the diffusion paths of two different datasets (CIFAR10 and SVHN) obtained via DDIM integration. The paths have different first and second derivatives (rate-of-change and curvature). We propose to measure these quantities for OOD detection.
@@
    Through comprehensive experiments with various OOD datasets, we show that a single generative model outperforms baselines that necessitates separate models for each distribution.
@@ 2 Background
${\mathrm{d}}{\mathbf{x}}_{t}=[{\bm{f}}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}({\mathbf{x}}_{t})]{\mathrm{d}}t+g(t)d\bar{{\mathbf{w}}}_{t},\ \ \ {\mathbf{x}}_{T}\sim p_{T}({\mathbf{x}})$ (2)
@@
We formalize the task of unsupervised OOD detection here. Given a distribution of interest $p({\mathbf{x}})$, the goal of OOD detection is to construct a scoring function which outputs a quantity $S_{\theta}({\mathbf{x}})\in\mathbb{R}$ that identifies if a given test point ${\mathbf{x}}_{\textrm{test}}$ is from $p({\mathbf{x}})$. In this work, a higher value of $S_{\theta}({\mathbf{x}}_{\textrm{test}})$ indicates that the sample is more likely to be drawn from $p({\mathbf{x}})$. We will use the notation “A vs B” to denote the task of distinguishing samples between A and B, where A is the in-distribution (ID) and B is the out-of-distribution (OOD) dataset. In unsupervised OOD detection, one must construct the function $S_{\theta}$ using only knowledge of A.
@@ 3 Diffusion Models for OOD Detection
An overview of our method, DiffPath, is illustrated in Fig. 1. DiffPath is based on the insight that the rate-of-change and curvature of the diffusion path connecting samples to standard normal differ between distributions, making them effective indicators for OOD detection. This section outlines the methodology behind DiffPath We begin in Sec. 3.1, where we provide evidence that likelihoods of a diffusion model are insufficient for OOD detection. Next, we show that the score function is a measure of the rate-of-change, as well as motivate the use of a single generative model in Sec. 3.2. Following which, we motivate the curvature as the derivative of the score in Sec. 3.3. We consider the curvature statistic as one variation of our method and abbreviate it as DiffPath-1D. In Sec. 3.4, we contextualize our method in terms of the optimal transport path between samples and standard normal, and finally propose a higher-order, hybrid variation called DiffPath-6D in Sec. 3.5, which incorporates both the rate-of-change and curvature quantities.
3.1 Likelihoods Do Not Work
@@
In Fig. 2, we plot the distributions of the negative ELBO (denoted NLL) of various training sets for a DM trained on CIFAR10. Our results corroborate earlier findings that likelihoods are not good OOD detectors; the NLL of CIFAR10 samples are higher than SVHN samples, meaning ID samples have lower likelihoods than OOD samples. Furthermore, the NLL of CIFAR10 overlaps significantly with that of CelebA samples. The poor AUROC score in Table 1 quantitatively demonstrates the inability of likelihoods to distinguish between ID and OOD samples. This motivates us to search for better statistics that we can extract from DMs for OOD detection.
3.2 Scores as OOD Statistic
@@
Method SVHN CelebA
NLL 0.091 0.573
$\sqrt{\sum_{t}\left\lVert\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\right\rVert_{2}^{2}}$ 0.856 0.638
$\sqrt{\sum_{t}\left\lVert\partial_{t}\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\right\rVert_{2}^{2}}$ 0.965 0.998
@@
The term $D_{\mathrm{KL}}(\phi_{T}\|\psi_{T})$ vanishes as $\phi_{T}=\psi_{T}=\mathcal{N}(\bm{0},{\mathbf{I}})$ by construction, assuming the true scores are available. In practice, one does not have access to the true score, and we rely on a score estimator $\bm{\epsilon}_{\theta}$ obtained via score matching approaches. Theorem 1 suggests that the scores of the marginal distributions along the ODE path serve as a proxy for the KL divergence: as $D_{\mathrm{KL}}(\phi_{0}\|\psi_{0})$ increases, so should the difference in the norms of their scores. Another interpretation is that this difference, $\mathbb{E}[\left\lVert\bm{\epsilon}_{\phi}({\mathbf{x}}_{t},t)-\bm{\epsilon}_{\psi}({\mathbf{x}}_{t},t)\right\rVert^{2}_{2}]$, is a measure of the Fisher divergence between the two distributions, which forms the foundation for score matching [17]. Thus, this motivates the use of the norm of the scores as an OOD statistic in distinguishing two distributions apart.
However, Theorem 1 is not immediately useful as it requires a priori knowledge of both distributions, whereas in unsupervised OOD detection only knowledge of the in-distribution is available. Interestingly, we empirically observe that it is possible to approximate the forward probability flow ODE for different distributions using a single diffusion model when it is trained on a large, diverse dataset $p_{0}({\mathbf{x}})$. Recall that as the PF ODE has the same marginal as the forward SDE, if the score estimate $\bm{\epsilon}_{\theta}$ has converged to the true score, then forward integration of a sample ${\mathbf{x}}_{0}$ using Eq. 6 should bring the sample to approximately standard normal, ${\mathbf{x}}_{T}\sim\mathcal{N}(\bm{0},{\mathbf{I}})$.
@@
where $\gamma_{t}=\sqrt{\frac{1-\bar{\alpha}_{t}^{2}}{\bar{\alpha}_{t}}}$ and $\bar{{\mathbf{x}}}_{t}={\mathbf{x}}_{t}\sqrt{1+\gamma_{t}^{2}}$. Let $\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)=-\sigma_{t}\nabla_{\mathbf{x}}\log p_{t}({\mathbf{x}})$ be a score model trained on $p_{0}({\mathbf{x}})$. It is known that the DDIM sampler [19] is Euler’s method applied to Eq. 7. In Fig. 3, we integrate Eq. 7 forward in time using DDIM for samples from distributions unseen by the model during training. We observe the surprising fact that when $p_{0}({\mathbf{x}})$ is ImageNet, the model is able to bring the samples approximately to the standard normal. However, this is not true for the model trained on SVHN, suggesting that $p_{0}$ needs to be a large and diverse dataset. We ablate the choice of $p_{0}$ in Sec. 5.2.
Figure 3: Illustration of the forward integration of Eq.  7 on samples from CIFAR10, SVHN and CelebA. The ImageNet model is able bring the samples to approximately standard normal despite never being trained explicitly on them. The SVHN model is only able to bring the SVHN sample to standard normal. This highlights the necessity for the diffusion model to be trained on a large, diverse dataset so that it can estimate the scores for a variety of data accurately.
This motivates $\bm{\epsilon}_{\theta}$ as a replacement for arbitrary $\bm{\epsilon}_{\phi}$ when integrating Eq. 7 forward with samples from $\phi_{0}$. In Fig. 2, we see that the distributions of $\sqrt{\sum_{t}\left\lVert\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\right\rVert_{2}^{2}}$ , the square root of the sum of $L^{2}$ norms of scores over time, applied to the three datasets using a single model trained on ImageNet are better separated than the likelihoods for CIFAR10 vs SVHN. We leverage this insight to design a simple scheme for OOD detection: fit a Kernel Density Estimator (KDE) to $\sqrt{\sum_{t}\left\lVert\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\right\rVert_{2}^{2}}$ of the training set for a given distribution, then use the KDE likelihoods of a test sample as the OOD score $S_{\theta}$. We provide pseudocode in Algorithm 1. The AUROC scores in Table 1 show a large improvement in CIFAR10 vs SVHN scores over likelihoods, although CIFAR10 vs CelebA is still lacking.
@@ 3.3 Second-Order Taylor Expansion (DiffPath-1D)
Surprisingly, we observe that for high-dimensional distributions such as images that we consider in this work, the time derivative in Eq. 10 alone provides an accurate enough estimate for OOD detection. Using the same ImageNet model as in Sec. 3.2, we see from Fig. 2 that the distributions of $\sqrt{\sum_{t}\left\lVert\partial_{t}\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\right\rVert_{2}^{2}}$ for different datasets are more spread out. This reflects quantitatively in the AUROC scores in Table 1, where the second-order statistic improves significantly upon the first-order scores, in particular for CIFAR10 vs CelebA with near perfect performance.
We thus consider the second-order statistic alone, $\sqrt{\sum_{t}\left\lVert\partial_{t}\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\right\rVert_{2}^{2}}$ , as one of our main statistic for OOD detection. As it is a one-dimensional quantity, we abbreviate it as DiffPath-1D.
@@ 3.5 Higher-dimensional Statistic (DiffPath-6D)
Owing to its simplicity, the one-dimensional statistic proposed in Sec. 3.3 may suffer from edge cases or perform suboptimally as information is condensed to a single scalar. For instance, given an image ${\mathbf{x}}_{0}$ with pixels normalized to the range $[-1,1]$, one such edge case is distinguishing ${\mathbf{x}}_{0}$ from itself with the sign of the pixels flipped, $-{\mathbf{x}}_{0}$. The two samples will produce symmetric OT paths differing only by a negative sign, resulting in identical statistics after taking the $L^{2}$ norm. We can see this from Table 2 where DiffPath-1D fails to distinguish CIFAR10 samples from itself with signs flipped, which we call negative CIFAR10. As such, we are motivated to propose a higher-dimensional statistic that does not utilize the standard form of the $L^{p}$ norm: $\left\lVert{\mathbf{x}}\right\rVert_{p}=\sum_{i}|{\mathbf{x}}_{i}|^{p}$. We define a new scalar quantity, $\langle{\mathbf{x}}\rangle_{p}=\sum_{i}({\mathbf{x}}_{i})^{p}$, which omits the absolute value so that sign information is retained, and propose a new six-dimensional statistic we dub DiffPath-6D:

### v2

@@ Abstract
Out-of-distribution (OOD) detection is a critical task in machine learning that seeks to identify abnormal samples. Traditionally, unsupervised methods utilize a deep generative model for OOD detection. However, such approaches require a new model to be trained for each inlier dataset. This paper explores whether a single model can perform OOD detection across diverse tasks. To that end, we introduce Diffusion Paths (DiffPath), which uses a single diffusion model originally trained to perform unconditional generation for OOD detection. We introduce a novel technique of measuring the rate-of-change and curvature of the diffusion paths connecting samples to the standard normal. Extensive experiments show that with a single model, DiffPath is competitive with prior work using individual models on a variety of OOD tasks involving different distributions. Our code is publicly available at https://github.com/clear-nus/diffpath.
@@ 1 Introduction
Out-of-distribution (OOD) detection, also known as anomaly or outlier detection, seeks to detect abnormal samples that are far from a given distribution. This is a vital problem as deep neural networks are known to be overconfident when making incorrect predictions on outlier samples [1, 2], leading to potential issues in safety-critical applications such as robotics, healthcare, finance, and criminal justice [3]. Traditionally, OOD detection using only unlabeled data relies on training a generative model on in-distribution data. Thereafter, measures such as model likelihood or its variants are used as an OOD detection score [4, 5, 6]. An alternative approach is to utilize the excellent sampling capabilities of diffusion models (DMs) to reconstruct corrupted samples, and use the reconstruction loss as an OOD measure [7, 8, 9].
However, these conventional methods require separate generative models tailored to specific inlier distributions and require retraining if the inlier data changes, such as in continual learning setups. This prompts the question: can OOD detection be performed using a single generative model? We answer in the affirmative and present Diffusion Paths (DiffPath) in this paper. While the use of a single model for OOD detection has been proposed in the discriminative setting [10], to the best of our knowledge, we are the first to explore this for generative models. We believe that the generative setting is particularly salient in light of recent trends where single generative foundation models are utilized across various tasks [11, 12].
Our method utilizes a single pretrained DM. In a departure from prior works that utilize variants of likelihoods [6, 5, 4] or reconstruction losses [7, 8, 9], we propose to perform OOD detection by measuring characteristics of the forward diffusion trajectory, specifically its rate-of-change and curvature, which can be computed from the score predicted by the diffusion model. We provide theoretical and empirical analyses that motivate these quantities as useful OOD detectors; their magnitudes are similar for samples from the same distribution and different otherwise. We summarize our contributions as follows:
Figure 1: Illustration of the diffusion paths of samples from two different distributions (CIFAR10 and SVHN) obtained via DDIM integration. The paths have different first and second derivatives (rate-of-change and curvature). We propose to measure these quantities for OOD detection.
@@
    Through comprehensive experiments with various datasets, we show that a single generative model is competitive with baselines that necessitates separate models for each distribution.
@@ 2 Background
${\mathrm{d}}{\mathbf{x}}_{t}=[{\bm{f}}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}({\mathbf{x}}_{t})]{\mathrm{d}}t+g(t){\mathrm{d}}\bar{{\mathbf{w}}}_{t},\ \ \ {\mathbf{x}}_{T}\sim p_{T}({\mathbf{x}})$ (2)
@@
Given a distribution of interest $p({\mathbf{x}})$, the goal of OOD detection is to construct a scoring function which outputs a quantity $S_{\theta}({\mathbf{x}})\in\mathbb{R}$ that identifies if a given test point ${\mathbf{x}}_{\textrm{test}}$ is from $p({\mathbf{x}})$. In this work, a higher value of $S_{\theta}({\mathbf{x}}_{\textrm{test}})$ indicates that the sample is more likely to be drawn from $p({\mathbf{x}})$. We will use the notation “A vs B” to denote the task of distinguishing samples between A and B, where A is the inlier distribution and B is the outlier distribution. In unsupervised OOD detection, one must construct the function $S_{\theta}$ using only knowledge of A.
@@ 3 Diffusion Models for OOD Detection
An overview of our method, DiffPath, is illustrated in Fig. 1. DiffPath is based on the insight that the rate-of-change and curvature of the diffusion path connecting samples to standard normal differ between distributions, making them effective indicators for OOD detection. This section outlines the methodology behind DiffPath. We begin in Sec. 3.1, where we provide evidence that likelihoods from a diffusion model are insufficient for OOD detection. Next, Sec. 3.2 shows that the score function is a measure of the rate-of-change and motivates the use of a single generative model. We then motivate the curvature as the derivative of the score in Sec. 3.3. We consider the curvature statistic as one variation of our method and abbreviate it as DiffPath-1D. In Sec. 3.4, we contextualize our method in terms of the optimal transport path between samples and standard normal, and finally propose a higher-order, hybrid variation called DiffPath-6D in Sec. 3.5, which incorporates both the rate-of-change and curvature quantities.
3.1 Diffusion Model Likelihoods Do Not Work for OOD Detection
@@
In Fig. 2, we plot the distributions of the negative ELBO (denoted NLL) of the CIFAR10 and SVHN training sets for a DM trained on CIFAR10. Our results corroborate earlier findings that likelihoods are not good OOD detectors; the NLL of CIFAR10 samples are higher than SVHN samples, meaning in-distribution samples have lower likelihoods than out-of-distribution samples. The poor AUROC score in Table 1 quantitatively demonstrates the inability of likelihoods to distinguish between inlier and outlier samples. This motivates us to search for better statistics that we can extract from DMs for OOD detection.
3.2 Scores as an OOD Statistic
@@
SVHN
NLL 0.091
$\sqrt{\sum_{t}\left\lVert\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\right\rVert_{2}^{2}}$ 0.856
$\sqrt{\sum_{t}\left\lVert\partial_{t}\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\right\rVert_{2}^{2}}$ 0.965
@@
The term $D_{\mathrm{KL}}(\phi_{T}\|\psi_{T})$ vanishes as $\phi_{T}=\psi_{T}=\mathcal{N}(\bm{0},{\mathbf{I}})$ by construction, assuming the true scores are available. In practice, we rely on a score estimator $\bm{\epsilon}_{\theta}$ obtained via score matching approaches. Theorem 1 suggests that the scores of the marginal distributions along the ODE path serve as a proxy for the KL divergence: as $D_{\mathrm{KL}}(\phi_{0}\|\psi_{0})$ increases, so should the difference in the norms of their scores. Another interpretation is that this difference, $\mathbb{E}[\left\lVert\bm{\epsilon}_{\phi}({\mathbf{x}}_{t},t)-\bm{\epsilon}_{\psi}({\mathbf{x}}_{t},t)\right\rVert^{2}_{2}]$, is a measure of the Fisher divergence between the two distributions, which forms the foundation for score matching [17]. This motivates using the norm of the scores as a statistic for distinguishing two distributions.
However, Theorem 1 is not immediately useful as it requires a priori knowledge of both distributions, whereas in unsupervised OOD detection, only knowledge of the inlier distribution is available. Interestingly, we empirically observe that it is possible to approximate the forward probability flow ODE for different distributions using a single diffusion model. Recall that as the PF ODE has the same marginal as the forward SDE, if the score estimate $\bm{\epsilon}_{\theta}$ has converged to the true score, then forward integration of a sample ${\mathbf{x}}_{0}$ using Eq. 6 should bring the sample to approximately standard normal, ${\mathbf{x}}_{T}\sim\mathcal{N}(\bm{0},{\mathbf{I}})$.
@@
where $\gamma_{t}=\sqrt{\frac{1-\bar{\alpha}_{t}^{2}}{\bar{\alpha}_{t}}}$ and $\bar{{\mathbf{x}}}_{t}={\mathbf{x}}_{t}\sqrt{1+\gamma_{t}^{2}}$. Let $\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)=-\sigma_{t}\nabla_{\mathbf{x}}\log p_{t}({\mathbf{x}})$ be a score model trained on $p_{0}({\mathbf{x}})$. It is known that the DDIM sampler [19] is Euler’s method applied to Eq. 7. In Fig. 3, we integrate Eq. 7 forward in time using DDIM for samples from various distributions, most of which are unseen by the model during training. Qualitatively, we observe the surprising fact that both the ImageNet and CelebA models are able to bring the samples approximately to the standard normal. We ablate the choice of $p_{0}$ in Sec. 5.2.
Figure 3: Illustration of the forward integration of Eq.  7 on samples from CIFAR10, SVHN and CelebA. Both the ImageNet and CelebA models are able bring the samples approximately to standard normal. Other than the case where the CelebA model is used to integrate CelebA samples (last row of the right figure), the samples shown here have not been seen by the models during training. While in certain cases the end result appears to contain features of the original image, thus deviating from an isotropic Gaussian (e.g., first row of the right figure), empirically we find that the scores remain accurate enough for outlier detection; see Sec. 5 for quantitative results.
@@
This motivates $\bm{\epsilon}_{\theta}$ as a replacement for arbitrary $\bm{\epsilon}_{\phi}$ when integrating Eq. 7 forward with samples from $\phi_{0}$. In Fig. 2, we see that the distributions of $\sqrt{\sum_{t}\left\lVert\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\right\rVert_{2}^{2}}$ , the square root of the sum of $L^{2}$ norms of scores over time, applied to the two datasets using a single model trained on ImageNet are better separated than the likelihoods. Note that Theorem 1 tells us only that the score norms of inlier and outlier samples are different, not whether one is higher or lower than the other. Thus, we propose the following OOD detection scheme: fit a Kernel Density Estimator (KDE) to $\sqrt{\sum_{t}\left\lVert\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\right\rVert_{2}^{2}}$ of the training set for a given distribution, then use the KDE likelihoods of a test sample as the OOD score $S_{\theta}$. This way, the likelihoods of outlier samples under the KDE are lower than for inlier samples, allowing us to compute the AUROC. We provide pseudocode in Algorithm 1. With this scheme, the AUROC scores in Table 1 show a large improvement over likelihoods; however, we would like to pursue even better OOD statistics, which we discuss in Sec. 3.3.
@@ 3.3 Second-Order Taylor Expansion (DiffPath-1D)
Surprisingly, we observe that for high-dimensional distributions such as images that we consider in this work, the time derivative in Eq. 10 alone provides an accurate enough estimate for OOD detection. Using the same ImageNet model as in Sec. 3.2, we observe an improvement in AUROC scores in CIFAR10 vs SVHN in Table 1 when using the second-order statistic. Qualitatively, the distributions of $\sqrt{\sum_{t}\left\lVert\partial_{t}\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\right\rVert_{2}^{2}}$ are more spread out in Fig. 2 as compared to the first-order scores, although the distinction is subtle; the quantitative results provide a more reliable confirmation of the benefit of using the second-order statistic.
We thus consider the second-order statistic alone, $\sqrt{\sum_{t}\left\lVert\partial_{t}\bm{\epsilon}_{\theta}({\mathbf{x}}_{t},t)\right\rVert_{2}^{2}}$ , as a possible statistic for OOD detection. As it is a one-dimensional quantity, we abbreviate it as DiffPath-1D. We evaluate DiffPath-1D in Sec. 5.2.
@@ 3.5 Higher-dimensional Statistic (DiffPath-6D)
Owing to its simplicity, the one-dimensional statistic proposed in Sec. 3.3 may suffer from edge cases or perform suboptimally as information is condensed to a single scalar. For instance, given an image ${\mathbf{x}}_{0}$ with pixels normalized to the range $[-1,1]$, one such edge case is distinguishing ${\mathbf{x}}_{0}$ from itself with the sign of the pixels flipped, $-{\mathbf{x}}_{0}$. The two samples will produce symmetric OT paths differing only by a negative sign, resulting in identical statistics after taking the $L^{2}$ norm. We can see this from Table 2 where DiffPath-1D fails to distinguish CIFAR10 samples from itself with signs flipped, which we call negative CIFAR10. As such, we propose a higher-dimensional statistic that does not utilize the standard form of the $L^{p}$ norm: $\left\lVert{\mathbf{x}}\right\rVert_{p}=\sum_{i}|{\mathbf{x}}_{i}|^{p}$. We define a new scalar quantity, $\langle{\mathbf{x}}\rangle_{p}=\sum_{i}({\mathbf{x}}_{i})^{p}$, which retains the sign information, and propose a new six-dimensional statistic we dub DiffPath-6D: